# Inference for Heterogeneous Treatment Effects with Efficient Instruments and Machine Learning

### v1

@@ D.4 Proof of Theorem 10
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=O_{P}(1)$ (69)
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}(\Delta\phi_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=O_{P}(1)$ (70)
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}(\Delta\phi_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=O_{P}(1)$ (71)
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{2}(\Delta\phi_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=O_{P}(1)$ (72)
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(\Delta\phi_{i}^{k})^{4}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=O_{P}(1).$ (73)
For (69), using the Cauchy-Schwarz inequality, it is enough to have $\frac{1}{nh}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{4}K\left(\frac{A_{i}-a}{h}\right)^{2}=O_{P}(1)$ and $\frac{1}{nh\sigma_{Q}^{4}(a)}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{4}K\left(\frac{A_{i}-a}{h}\right)^{2}=O_{P}(1)$. The expectation of the first term is bounded by assertion 14 of Assumption 6 and hence the term is $O_{P}(1)$ by Lemma 26. The second follows by taking the expectation conditional on $I_{k}^{c}$ since $\|f-\hat{f}\|_{L_{4}}/\sigma_{Q}(a)=O_{P}\left(h^{1/4}\right)$ by the assumption of Theorem 10. Hence, (69) follows and (70) follows in exactly the same way. Finally, (71), (72) and (73) follow from (49), (53) and (52). This concludes the proof of (68) and hence also the proof of Theorem 10.
@@ D.5 Some Lemmas
Assume that the kernel $K$ satisfies Assumption 3. Moreover, assume that the function $\rho:\mathbb{R}\to\mathbb{R}$ defined by $\rho(x)=v(x)\cdot p_{A}(x)$ for some function $v(\cdot)$ is two times differentiable for $x$ in the support of $A_{i}$. Then,
@@
    $\mathbb{E}\left[K\left(\frac{A_{i}-a}{h}\right)v(A_{i})\right]=h\rho(a)+h^{3}\frac{\rho^{\prime\prime}(a)}{2}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x+R(a,h)$
@@
    $|R(a,h)|\leq h^{3}\Biggl{(}\|\rho^{\prime\prime}\|_{\infty}\int_{|x|>h^{-1/2}}K(x)x^{2}\mathrm{d}x$
    $\qquad\qquad\qquad\quad+\frac{1}{2}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x\cdot\sup_{|s-a|\leq h^{1/2}}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(a)|\Biggr{)}.$
    In particular, if $\rho^{\prime\prime}$ is continuous at $a$, then $R(a,h)=o(h^{3})$. If $|\rho^{\prime\prime}|$ bounded but not continuous at $a$, we still have $R(a,h)=O(h^{3})$.
@@
    $\mathbb{E}\left[K\left(\frac{A_{i}-a}{h}\right)^{2}v(A_{i})\right]$
    $=h\rho(a)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x+h^{3}\frac{\rho^{\prime\prime}(a)}{2}\int_{-\infty}^{\infty}K(x)^{2}x^{2}\mathrm{d}x+\bar{R}(a,h)$
@@
    $\qquad\qquad\qquad\quad+\frac{1}{2}\int_{-\infty}^{\infty}K(x)^{2}x^{2}\mathrm{d}x\cdot\sup_{|s-a|\leq h^{1/2}}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(a)|\Biggr{)}.$
    In particular, if $\rho^{\prime\prime}$ is continuous at $a$, then $\bar{R}(a,h)=o(h^{3})$. If $|\rho^{\prime\prime}|$ bounded but not continuous at $a$, we still have $\bar{R}(a,h)=O(h^{3})$.
@@
We only prove the first statement. The second statement follows by replacing $K$ with $K^{2}$ in the proof. From Taylor’s theorem, we can write for some $\xi_{x}\in[a,x]$ (or in $[x,a]$)
$\rho(x)=\rho(a)+\rho^{\prime}(a)(x-a)+\rho^{\prime\prime}(a)\frac{(x-a)^{2}}{2}+(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(a))\frac{(x-a)^{2}}{2}.$
@@
$\mathbb{E}\left[K\left(\frac{A_{i}-a}{h}\right)v(A_{i})\right]$ $=\int K\left(\frac{x-a}{h}\right)v(x)p_{A}(x)\mathrm{d}x$
$=\int K\left(\frac{x-a}{h}\right)\rho(x)\mathrm{d}x$
$=\rho(a)\int K\left(\frac{x-a}{h}\right)\mathrm{d}x$
$+\frac{\rho^{\prime\prime}(a)}{2}\int K\left(\frac{x-a}{h}\right)(x-a)^{2}\mathrm{d}x+R(a,h)$
$=h\rho(a)+h^{3}\frac{\rho^{\prime\prime}(a)}{2}\int K(x)x^{2}\mathrm{d}x+R(a,h)$
@@
$R(a,h)=\int K\left(\frac{x-a}{h}\right)\left(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(a)\right)\frac{(x-a)^{2}}{2}\mathrm{d}x.$
@@
$|R(a,h)|$
$\leq|\int_{|x-a|>th}K\left(\frac{x-a}{h}\right)\left(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(a)\right)\frac{(x-a)^{2}}{2}\mathrm{d}x|$
$+|\int_{|x-a|\leq th}K\left(\frac{x-a}{h}\right)\left(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(a)\right)\frac{(x-a)^{2}}{2}\mathrm{d}x|$
$\leq\int_{|x-a|>th}K\left(\frac{x-a}{h}\right)2\|\rho^{\prime\prime}\|_{\infty}\frac{(x-a)^{2}}{2}\mathrm{d}x$
$+\int_{|x-a|\leq th}K\left(\frac{x-a}{h}\right)\sup_{|s-a|\leq ht}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(a)|\frac{(x-a)^{2}}{2}\mathrm{d}x$
$\leq h^{3}\left(\|\rho^{\prime\prime}\|_{\infty}\int_{|x|>t}K(x)x^{2}\mathrm{d}x+\frac{1}{2}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x\cdot\sup_{|s-a|\leq ht}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(a)|\right).$
@@ Appendix E Further Simulations
E.1 Multidimensional $X$
We keep the framework of Section 4.2 but consider $5$-dimensional covariates $X_{i}$. We use the data generating mechanism
$X_{i}$ $\sim_{i.i.d.}\mathcal{N}_{5}(0,\Sigma),\,\Sigma_{i,j}=0.5+0.5\cdot 1\{i=j\},$ (74)
@@ E.1 Multidimensional $X$
$\delta_{i}$ $\leftarrow 0.7H_{i}+0.7E_{\delta,i},$
@@
$A_{i}$ $\leftarrow X_{i,1},$
$Y_{i}$ $\leftarrow\beta(A_{i})D_{i}+g(X_{i})+\epsilon_{i}.$
Note that the components of $X_{i}$ all have variance $1$ and are correlated. Moreover, we change the function $g$ to $g(x)=\tanh(x_{1})-x_{3}$ and change the settings (Z lin.) and (Z nonlin.) to
@@
Figure 6: Simulation results for setting (hom.)/(Z lin.) (top) and (hom.)/(Z nonlin.) (bottom) for 5-dimensional $X_{i}$. Left panel: mean squared error (MSE) of $\hat{\beta}$ (hom. T.E.), $\hat{\beta}(0)$ (het. T.E. for $a=0$)and $\hat{\beta}(1.5)$ (het. T.E. for $a=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Solid lines correspond to the methods (hom. T.E. linearIV) and (hom. T.E. mlIV). Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $a=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $a=1.5$.
@@
The results are very similar to before: in the (Z lin.) settings, the estimators based on (linearIV) exhibit almost identical results to the estimators based on (mlIV), whereas in the (Z nonlin.) settings, the estimators based on (mlIV) have significantly smaller MSEs, while still exhibiting good empirical coverage of both standard and robust confidence intervals.
Secondly, we consider the settings (het.)/(Z lin.) and (het.)/(Z nonlin.) and use the methods (het. T.E. mlIV) and (het. T.E. linearIV) for $\hat{\beta}(0)$ and $\hat{\beta}(1.5)$. The results are shown in Figure 7. Again, the results are comparable to those from Section 4.2. There is one surprising exception however: the coverage of the robust confidence interval for (het. T.E. mlIV) for $\hat{\beta}(1.5)$ is not good for large values of $N$. Note that however, the confidence intervals are only guaranteed to be asymptotically valid if we can consistently estimate the nuisance functions at sufficiently fast rates. However, from the data generating process (74) in setting (het.)/(Z nonlin.), we see that $\mathbb{E}[D_{i}|X_{i}]$ and $\mathbb{E}[Z_{i}|X_{i}]$ are not additive functions in the components of $X_{i}$, so gam will in general not yield a consistent estimate of all the nuisance functions.
@@
Figure 7: Simulation results for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom) for 5-dimensional $X_{i}$. Left panel: mean squared error (MSE) of $\hat{\beta}(0)$ (het. T.E. for $a=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $a=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $a=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $a=1.5$.
To get further insights into this issue, consider Figure 8, where we consider exactly the same simulation scenarios (het.)/(Z lin.) and (het.)/(Z nonlin.), but instead of gam, we use boosted regression trees using the R package xgboost [9, 10].
Figure 8: Same as Figure 7, but the nuisance functions are estimated using xgboost instead of gam.
Now, we see that also the coverage of the robust confidence set in the setting (het.)/(Z nonlin.) is around 0.95. We attribute this to the fact that xgboost can – in contrast to the additive model fitted by gam – approximate a function space that is rich enough to capture all the conditional expectations needed. The other plots look similar to before.

### v2

@@ D.4 Proof of Theorem 9
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=O_{P}(1)$ (69)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=O_{P}(1)$ (70)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}(\Delta\phi_{1,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=O_{P}(1)$ (71)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{2}(\Delta\phi_{1,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=O_{P}(1)$ (72)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta\phi_{1,i}^{k})^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=O_{P}(1).$ (73)
For (69), using the Cauchy-Schwarz inequality, it is enough to have $\frac{1}{nh}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{4}K\left(\frac{V_{i}-v}{h}\right)^{2}=O_{P}(1)$ and $\frac{1}{nh\sigma_{Q}^{4}(v)}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{4}K\left(\frac{V_{i}-v}{h}\right)^{2}=O_{P}(1)$. The expectation of the first term is bounded by assertion 14 of Assumption 6 and hence the term is $O_{P}(1)$ by Lemma 25. The second follows by taking the expectation conditional on $I_{k}^{c}$ since $\|f-\hat{f}\|_{L_{4}}/\sigma_{Q}(v)=O_{P}\left(h^{1/4}\right)$ by the assumption of Theorem 9. Hence, (69) follows and (70) follows in exactly the same way. Finally, (71), (72) and (73) follow from (49), (53) and (52). This concludes the proof of (68) and hence also the proof of Theorem 9.
@@ D.5 Some Lemmas
Assume that the kernel $K$ satisfies Assumption 3. Moreover, assume that the function $\rho:\mathbb{R}\to\mathbb{R}$ defined by $\rho(x)=v(x)\cdot p_{V}(x)$ for some function $v(\cdot)$ is two times differentiable for $x$ in the support of $V_{i}$. Then,
@@
    $\mathbb{E}\left[K\left(\frac{V_{i}-v}{h}\right)v(V_{i})\right]=h\rho(v)+h^{3}\frac{\rho^{\prime\prime}(v)}{2}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x+R(v,h)$
@@
    $|R(v,h)|\leq h^{3}\Biggl(\|\rho^{\prime\prime}\|_{\infty}\int_{|x|>h^{-1/2}}K(x)x^{2}\mathrm{d}x$
    $\qquad\qquad\qquad\quad+\frac{1}{2}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x\cdot\sup_{|s-v|\leq h^{1/2}}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(v)|\Biggr).$
    In particular, if $\rho^{\prime\prime}$ is continuous at $v$, then $R(v,h)=o(h^{3})$. If $|\rho^{\prime\prime}|$ bounded but not continuous at $v$, we still have $R(v,h)=O(h^{3})$.
@@
    $\mathbb{E}\left[K\left(\frac{V_{i}-v}{h}\right)^{2}v(V_{i})\right]$
    $=h\rho(v)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x+h^{3}\frac{\rho^{\prime\prime}(v)}{2}\int_{-\infty}^{\infty}K(x)^{2}x^{2}\mathrm{d}x+\bar{R}(v,h)$
@@
    $\qquad\qquad\qquad\quad+\frac{1}{2}\int_{-\infty}^{\infty}K(x)^{2}x^{2}\mathrm{d}x\cdot\sup_{|s-v|\leq h^{1/2}}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(v)|\Biggr).$
    In particular, if $\rho^{\prime\prime}$ is continuous at $v$, then $\bar{R}(v,h)=o(h^{3})$. If $|\rho^{\prime\prime}|$ bounded but not continuous at $v$, we still have $\bar{R}(v,h)=O(h^{3})$.
@@
We only prove the first statement. The second statement follows by replacing $K$ with $K^{2}$ in the proof. From Taylor’s theorem, we can write for some $\xi_{x}\in[v,x]$ (or in $[x,v]$)
$\rho(x)=\rho(v)+\rho^{\prime}(v)(x-v)+\rho^{\prime\prime}(v)\frac{(x-v)^{2}}{2}+(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(v))\frac{(x-v)^{2}}{2}.$
@@
$\mathbb{E}\left[K\left(\frac{V_{i}-v}{h}\right)v(V_{i})\right]$ $=\int K\left(\frac{x-v}{h}\right)v(x)p_{V}(x)\mathrm{d}x$
$=\int K\left(\frac{x-v}{h}\right)\rho(x)\mathrm{d}x$
$=\rho(v)\int K\left(\frac{x-v}{h}\right)\mathrm{d}x$
$+\frac{\rho^{\prime\prime}(v)}{2}\int K\left(\frac{x-v}{h}\right)(x-v)^{2}\mathrm{d}x+R(v,h)$
$=h\rho(v)+h^{3}\frac{\rho^{\prime\prime}(v)}{2}\int K(x)x^{2}\mathrm{d}x+R(v,h)$
@@
$R(v,h)=\int K\left(\frac{x-v}{h}\right)\left(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(v)\right)\frac{(x-v)^{2}}{2}\mathrm{d}x.$
@@
$|R(v,h)|$
$\leq|\int_{|x-v|>th}K\left(\frac{x-v}{h}\right)\left(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(v)\right)\frac{(x-v)^{2}}{2}\mathrm{d}x|$
$+|\int_{|x-v|\leq th}K\left(\frac{x-v}{h}\right)\left(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(v)\right)\frac{(x-v)^{2}}{2}\mathrm{d}x|$
$\leq\int_{|x-v|>th}K\left(\frac{x-v}{h}\right)2\|\rho^{\prime\prime}\|_{\infty}\frac{(x-v)^{2}}{2}\mathrm{d}x$
$+\int_{|x-v|\leq th}K\left(\frac{x-v}{h}\right)\sup_{|s-a|\leq ht}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(v)|\frac{(x-v)^{2}}{2}\mathrm{d}x$
$\leq h^{3}\left(\|\rho^{\prime\prime}\|_{\infty}\int_{|x|>t}K(x)x^{2}\mathrm{d}x+\frac{1}{2}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x\cdot\sup_{|s-v|\leq th}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(v)|\right).$
@@ Appendix E Further Simulations
E.1 Varying IV Strength
We consider the same simulation setup (26) as in Section 4.2 with two differences:
  1. 1.
    We introduce an additional parameter $s$ (for the IV strength) to the function $f(z,x)$, namely
    (Z lin.)
    $f(z,x)=-\sin(x)+s\cdot z$,
    (Z nonlin.)
    $f(z,x)=-\sin(x)+s\cdot(\cos(z)+0.2z)$
    If the strength $s=0$, the treatment $D_{i}$ is not related to the instrument $Z_{i}$. If $s$ increases, the strength of the instrument $Z_{i}$ increases.
  2. 2.
    We replace the definitions of $\delta_{i}$ and $\epsilon_{i}$ by
    $\delta_{i}$ $\leftarrow 0.7H_{i}+0.1E_{\delta,i},$
    $\epsilon_{i}$ $\leftarrow 0.7H_{i}+0.1E_{\epsilon,i}$
    to make them more correlated, i.e., increase the amount of endogeneity.
We fix $N=500$ and vary the strength $s\in\{0,0.1,\ldots,1\}$. Again, we simulate $n_{\text{sim}}=1000$ datasets and use generalized additive models (gam) for the nuisance functions, the Epanechnikov kernel, and choose the bandwidth as the normal reference rule times $N^{1/5}/N^{1/7}$. The corresponding plots for the settings (hom.)/(Z lin.) and (hom.)/(Z nonlin.) can be found in Figure 6 and the plots for the settings (het.)/(Z lin.) and (het.)/(Z nonlin.) can be found in Figure 7.
Figure 6: Simulation results for setting (hom.)/(Z lin.) (top) and (hom.)/(Z nonlin.) (bottom) with varying IV strength $s$. Left panel: mean squared error (MSE) of $\hat{\beta}$ (hom. T.E.), $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Solid lines correspond to the methods (hom. T.E. linearIV) and (hom. T.E. mlIV). Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
Figure 7: Simulation results for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom) with varying IV strength $s$. Left panel: mean squared error (MSE) of $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
In both figures, we see that the mean squared error tends to decrease with increasing IV strength. Moreover, for small to moderately small IV strengths, the standard confidence intervals suffer from severe undercoverage and for the (Z nonlin.) settings, the undercoverage is worse for the methods based on (linearIV). In contrast, the robust confidence sets exhibit empirical coverage close to the nominal level of 0.95 regardless of the IV strength.
E.2 Multidimensional $X$
We keep the framework of Section 4.2 but use 5-dimensional covariates $X_{i}$ and use boosted regression trees using the R package xgboost [12, 13] for the nuisance function estimations. We slightly alter the simulation setup (26) and generate data according to
$X_{i}$ $\sim_{i.i.d.}\mathcal{N}_{5}(0,\Sigma),\,\Sigma_{i,j}=0.5+0.5\cdot 1\{i=j\},$
@@ E.2 Multidimensional $X$
$\delta_{i}$ $\leftarrow 0.7H_{i}+0.7E_{\delta,i},$ (74)
@@
$V_{i}$ $\leftarrow X_{i,1},$
$Y_{i}$ $\leftarrow\beta(V_{i})D_{i}+g(X_{i})+\epsilon_{i}.$
Note that the components of $X_{i}$ all have variance 11 and are correlated. Moreover, we change the function $g$ to $g(x)=\tanh(x_{1})-x_{3}$ and change the settings (Z lin.) and (Z nonlin.) to
@@
Figure 8: Simulation results for setting (hom.)/(Z lin.) (top) and (hom.)/(Z nonlin.) (bottom) for 5-dimensional $X_{i}$ and nuisance function estimation using xgboost. Left panel: mean squared error (MSE) of $\hat{\beta}$ (hom. T.E.), $\hat{\beta}(0)$ (het. T.E. for $v=0$)and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Solid lines correspond to the methods (hom. T.E. linearIV) and (hom. T.E. mlIV). Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
@@
The results are very similar to Section 4.2: in the (Z lin.) settings, the estimators based on (linearIV) exhibit almost identical results to the estimators based on (mlIV), whereas in the (Z nonlin.) settings, the estimators based on (mlIV) have significantly smaller MSEs, while still exhibiting good empirical coverage of both standard and robust confidence intervals.
Secondly, we consider the settings (het.)/(Z lin.) and (het.)/(Z nonlin.) and use the methods (het. T.E. mlIV) and (het. T.E. linearIV) for $\hat{\beta}(0)$ and $\hat{\beta}(1.5)$. The results are shown in Figure 9. Again, the results are comparable to those from Section 4.2.
@@
Figure 9: Simulation results for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom) for 5-dimensional $X_{i}$ and nuisance function estimation using xgboost. Left panel: mean squared error (MSE) of $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
Finally, let us consider the influence of the IV strength in the setting with 5-dimensional $X_{i}$. As in Section E.1, we include a parameter $s$ corresponding to different IV strengths and modify $\delta_{i}$ and $\epsilon_{i}$ to increase their correlation. Otherwise, we keep the setup from (74). The resulting plots for the settings (hom.)/(Z lin.) and (hom.)/(Z nonlin.) can be found in Figure 10 and the plots for the settings (het.)/(Z lin.) and (het.)/(Z nonlin.) can be found in Figure 11.
Figure 10: Simulation results for setting (hom.)/(Z lin.) (top) and (hom.)/(Z nonlin.) (bottom) with varying IV strength $s$, 5-dimensional $X_{i}$ and nuisance function estimation using xgboost. Left panel: mean squared error (MSE) of $\hat{\beta}$ (hom. T.E.), $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Solid lines correspond to the methods (hom. T.E. linearIV) and (hom. T.E. mlIV). Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
Figure 11: Simulation results for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom) with varying IV strength $s$, 5-dimensional $X_{i}$ and nuisance function estimation using xgboost. Left panel: mean squared error (MSE) of $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
Whereas for the (hom.) settings, there is still a considerable (though less pronounced) undercoverage of the standard confidence intervals for small IV strength, this effect seems to be almost gone for the (het.) settings.